Solve
5/4x+1/8x=3/8+x
This is how I did on my own
5/4x+1/8x-3/8-x=0
10/8x+1/8x-3/8-8/8=0
10x+x = 11x
11/8x-11/8=0
0=0
The answer is 1
How did they get this?
jpa

Answer by josmiceli(19441)

Solve
5/4x+1/8x=3/8+x
This is how I did on my own
5/4x+1/8x-3/8-x=0
-------------------------------------
10/8x+1/8x-3/8-8/8=0 this line should be:
10/8x + 1/8x - 3/8 - (8/8)x = 0
-------------------------------------
3/8x = 3/8
x = 1
--------
10x+x = 11x
11/8x-11/8=0
0=0
The answer is 1
How did they get this?
jpa